### v1

@@ Abstract
The Gaussian process (GP) is a widely used probabilistic machine learning method for stochastic function approximation, stochastic modeling, and analyzing real-world measurements of nonlinear processes. Unlike many other machine learning methods, GPs include an implicit characterization of uncertainty, making them extremely useful across many areas of science, technology, and engineering. Traditional implementations of GPs involve stationary kernels (also termed covariance functions) that limit their flexibility and exact methods for inference that prevent application to data sets with more than about ten thousand points. Modern approaches to address stationarity assumptions generally fail to accommodate large data sets, while all attempts to address scalability focus on approximating the Gaussian likelihood, which can involve subjectivity and lead to inaccuracies. In this work, we explicitly derive an alternative kernel that can discover and encode both sparsity and nonstationarity. We embed the kernel within a fully Bayesian GP model and leverage high-performance computing resources to enable the analysis of massive data sets. We demonstrate the favorable performance of our novel kernel relative to existing exact and approximate GP methods across a variety of synthetic data examples. Furthermore, we conduct space-time prediction based on more than one million measurements of daily maximum temperature and verify that our results outperform state-of-the-art methods in the Earth sciences. More broadly, having access to exact GPs that use ultra-scalable, sparsity-discovering, nonstationary kernels allows GP methods to truly compete with a wide variety of machine learning methods.
@@ 1 Introduction
Gaussian processes (GPs) are a preeminent framework for stochastic function approximation, statistical modeling of real-world measurements, and non-parametric and nonlinear regression within machine learning (ML) and surrogate modeling. One reason for the popularity of GPs is that they are analytically tractable and can be fully specified in terms of a prior mean function and a kernel, also known as a covariance function. Additionally, unlike most ML methods, GPs are “probabilistic” in the sense that they naturally include measures of uncertainty in predictions and, when implemented in a Bayesian setting, can furthermore account for and propagate uncertainty in their inferred hyperparameters. GPs have been broadly used across many areas of science, technology, and engineering, including forward modeling [1, 2, 3], uncertainty quantification [4, 5], and autonomous experimentation [6, 7, 8, 9].
@@
However, there are two main reasons why GPs are not used more broadly for analyzing modern data sets. First, traditional implementations of GPs involve stationary kernels [e.g., the squared exponential kernel is widely used in Gaussian process applications; 10] that are too inflexible to model real-world data [11]. The fact that standard stationary kernels are so widely used leads to the general impression that GPs themselves are inflexible, as opposed to a particular implementation of GPs being inflexible. Second, evaluating the Gaussian likelihood for $N$ data points scales with $\mathcal{O}(N^{2})$ in memory and $\mathcal{O}(N^{3})$ in computation, such that GPs are computationally infeasible for as few as several thousand data points. Given the fact that modern data sets quite commonly involve millions or even billions of data points, in an ML context, this lack of scalability leads users to gravitate towards neural networks and deep learning methods that are easily applied to big data.
The statistics and ML literature contains a wide variety of methods to address these two primary limitations of GPs. Much work has been done to derive and implement nonstationary kernels, wherein the dependence between two input locations cannot be written as a function of their separation distance [see, e.g., 12, 13, 14, 15, 16, 17]. However, with a few exceptions, these methods are generally difficult to implement for large data sets due to their large numbers of hyperparameters and the need for specialized algorithms for training. These methods can also lead to overfitting; multiple authors [18, 19, 20] have pointed out the potential dangers of using over-parameterized models (such as GPs with nonstationary kernels) without regularization techniques such as posterior averaging [21] or prior regularization [22, 23] in Bayesian non-parametric models [24, 3, 25].
@@
Regarding scalability, a variety of methods have been proposed that derive approximations to the GP likelihood, e.g., local GP experts [26], inducing-point and low-rank methods [27, 28, 29], structure-exploiting methods [30], and Vecchia approximations [31, 32, 33]. We refer the interested reader to [34] for a direct comparison of many of these methods on a common data set. These so-called “approximate GP” methods are not, in general, agnostic to the underlying kernel, and theoretical properties are, in most cases, derived only for stationary kernels. Furthermore, each of these methods requires multiple subjective choices – e.g., Vecchia approximations necessitate specifying the number of neighbors to condition on, the ordering of points, and the selection of nearest neighbors – that may introduce sensitive ad hoc choices in a given analysis.
@@
A final limitation of most kernels used in the literature (both stationary and nonstationary) is that they cannot model covariances that are exactly zero. In almost all cases, a kernel will decay to zero as the separation between two points increases; however, the kernel will still return covariances that are technically nonzero, even for points that are very far apart in the input space. It is our hypothesis that most (if not all) real-world data sets implicitly include at least some degree of sparsity in the covariance, but typical classes of kernels cannot flexibly model both zero and nonzero covariances. Compactly-supported kernels [see, e.g., 35, 36, 37, 38, 39] can model both zero and nonzero covariances in a data-driven manner, albeit usually with stationary kernels. The approximate GP methods also “model” sparsity, but the implied sparsity is fully determined by ad hoc choices such as the ordering of points and neighbor selection. The combination of nonstationary and also compactly supported kernels would enable more appropriate modeling of real-world data and yield computational shortcuts for the matrix operations needed to calculate the Gaussian likelihood. In short, the density of the GP prior covariance is not fundamental to the method itself but a result of poorly chosen kernels. Instead of using these standard kernels and then retroactively working on scalability, we propose to give the GP the ability to discover sparsity during the model fitting process. A demonstration of our proposed methodology is shown in Figure 1.
@@
Figure 1: Ground truth function over $[0,10]$ with 50 noisy measurements (panel a.); the synthetic function is both sparse and nonstationary. Panel (b) shows the resulting posterior mean function with 95% uncertainty intervals for four methods: a stationary Gaussian process with a Matérn (non-sparse) kernel, the Vecchia approximation (a leading approximate GP method), and our sparsity-discovering kernel combined with both stationary and nonstationary core kernels. Panel (c) shows posterior median covariance estimates (gray areas indicate where the kernel is identically zero). Our kernel has three clear benefits: first, it can discover sparsity in the data while using an exact Gaussian likelihood; second, it can infer the nonstationary structure in the data wherein diagonal elements of the covariance are larger at the endpoints of the domain; third, it outperforms exact and leading approximate methods in terms of root mean square error (RMSE) and continuous rank probability score (CRPS; lower scores for both metrics indicate a better fit).
@@
In this paper, we introduce a novel kernel that explicitly demonstrates how to apply the proof-of-concept methodology proposed in [40] to real-world data sets. The overall kernel consists of a compactly supported kernel that can discover and encode sparsity combined with an arbitrary “core” kernel, yielding ultra-flexible kernels that allow one to apply exact nonstationary GPs to arbitrarily large data sets. Together with prior regularization, the induced sparsity also prevents overfitting and leads to a better interpretation of the data since it preserves the flexibility of a nonstationary kernel. We also propose a complete Bayesian stochastic model based on the new kernel that accounts for a generic prior mean function and measurement error and allows posterior prediction for unobserved input locations with uncertainty quantification. While high-performance computing (HPC) resources are needed to apply our method to truly massive data sets, we also show how the proposed kernel can be used to conduct fully Bayesian analysis for moderately sized data sets with limited computational resources, e.g., using a personal laptop. The efficacy of our novel kernel is demonstrated using both synthetic data examples and a real-world data set from the Earth sciences.
@@ 2 Sparsity-discovering nonstationary kernels
where $C_{\text{core}}$ is a positive definite kernel (stationary or nonstationary), $C_{\text{sparse}}$ is a sparsity-inducing kernel (defined below in Eq. 2), and ${\boldsymbol{\theta}}_{y}=({\boldsymbol{\theta}}_{\text{core}},{\boldsymbol{\theta}}_{\text{sparse}})$. So long as $C_{\text{core}}$ and $C_{\text{sparse}}$ are positive definite, their product will be positive definite [41], yielding a powerful framework for building flexible kernels.
@@ 2.1 Nonstationary kernels that discover sparsity
$g({\bf x})=\left\{\begin{array}[]{ll}a\exp\left\{b\left[1-(1-||\mathbf{x}-\mathbf{h}||^{2}/r^{2})^{-1}\right]\right\}&\text{if }||\mathbf{x}-\mathbf{h}||^{2}<r\\ 0&\text{else},\end{array}\right.$ (1)
where $a\geq 0$ represents the amplitude, $b>0$ is a shape parameter (see Figure G.1 for a visualization of the effect of $b$), $\mathbf{h}\in\mathbb{R}^{d}$ is the centroid, and $r>0$ is the radius [42]. Note that $g({\bf x})$ is smooth and exactly zero for point-to-centroid distances greater than or equal to $r$, which is what enables the covariance matrix to be sparse. We combine these bump functions to yield a sparse kernel
@@
$f_{i}({\bf x})=\sum_{j=1}^{n_{2}}g_{ij}({\bf x}),\hskip 17.22217pti=1,\dots,n_{1},$
$f_{0}({\bf x},{\bf x}^{\prime};r_{0})$ is a compactly supported stationary kernel, and $s_{0}$ is a scalar parameter to adjust the relative magnitude of $f_{0}({\bf x},{\bf x};r_{0})$ and the amplitudes $\{a_{ij}\}$. We include an additive stationary kernel in Eq. 2 because there might otherwise be cases where $C({\bf x},{\bf x};{\boldsymbol{\theta}}_{\text{sparse}})=0$; specifying the stationary kernel to be compactly supported ensures that the resulting covariance matrix is still sparse. Among the broad range of possibilities for specifying a compactly supported $f_{0}$ [see, e.g., 36, 37, 38, 39], we opt to use members of the Wendland family of polynomials [35], e.g.,
$f_{0}({\bf x},{\bf x}^{\prime};r_{0})=\left\{\begin{array}[]{ll}\left(1-\frac{||{\bf x}-{\bf x}^{\prime}||}{r_{0}}\right)^{8}\left(35\left(\frac{||{\bf x}-{\bf x}^{\prime}||}{r_{0}}\right)^{3}+25\left(\frac{||{\bf x}-{\bf x}^{\prime}||}{r_{0}}\right)^{2}+\frac{8||{\bf x}-{\bf x}^{\prime}||}{r_{0}}+1\right)&\text{if }||{\bf x}-{\bf x}^{\prime}||^{2}<r_{0},\\ 0&\text{else.}\end{array}\right.$
@@
In order to show how the various components of $C_{\text{sparse}}$ in Eq. 2 combine to flexibly model sparsity in the overall kernel $C_{y}$, we present an illustrative example. For this example, we set $n_{1}=2$ and $n_{2}=3$ and take $\mathcal{X}=[0,1]$ with $N=100$ equally spaced locations (hence $d=1$). Figure 2(a) shows the bump functions $g_{ij}(x)$ and their sums $f_{i}(x)$ using fixed values of the bump function hyperparameters. Starting with the leftmost column of Figure 2, the product of the sums-of-bump-functions $\sum_{i}f_{i}(x)f_{i}(x^{\prime})$ in the middle row reveals the effect of the bump functions: for a given $i$, the support (i.e., domain of nonzero values) of the bump functions define distance-unrelated parts of the domain that are correlated; furthermore, across the various $i$, a given point in the input space can have a nonzero correlation with different sets of points that does not depending on the distances from this point to these sets. For example, note that both because $f_{1}(0.17)>0$ and $f_{1}(0.7)>0$, $C_{\text{sparse}}(0.17,0.7)>0$ even though these points are relatively far apart. Furthermore, because $f_{2}(0.17)>0$ and $f_{2}(0.9)>0$ (but $f_{2}(0.7)=0$), we have $C_{\text{sparse}}(0.17,0.9)>0$ but $C_{\text{sparse}}(0.7,0.9)=0$. While the radii $r_{ij}$ specify regions of zero versus nonzero correlation, the amplitudes $a_{ij}$ and shapes $b_{ij}$ determine the relative magnitude of off-diagonal nonzero correlations and how sharply peaked the correlations are near the bump function centers, respectively shown in Figure 2(b) and (c). Finally, increasing the radii in Figure 2(d) predictably decreases the overall sparsity of $C_{\text{sparse}}$ and (together with relatively small $b$) increases the areas of high correlations. The compactly supported stationary covariance $f_{0}(x,x^{\prime};r_{0}=0.15)$ serves to ensure that the diagonal elements of $C_{\text{sparse}}$ are nonzero, with $r_{0}$ determining the extent to which points very close together are correlated within the sparse kernel.
(a) Baseline example         (b) Shape×$\times 0.1$             (c) Amplitude×$\times 2$             (d) Radius×$\times 1.5$
@@
Our specification of $C_{\text{sparse}}$ implies a nonstationary kernel, irrespective of the choice of $C_{\text{core}}$: note that we cannot write Eq. 2 as a function of the separating vector ${\bf x}-{\bf x}^{\prime}$ since the bump functions in Eq. 1 cannot be written as a function of ${\bf x}-{\bf x}^{\prime}$. Our approach explicitly models sparsity directly in the covariance matrix to capture the inherent sparsity of the dependence structure in real-world datasets; alternatively, approximate methods [e.g., inducing points; 43, 29] focus on approximating the GP with a subset of the data to achieve computational savings. In other words, both classes of methods leverage sparsity to achieve computational savings, but the sparsity implied by our method is data driven and describes the underlying structure in the data set of interest.

### v3

@@ Abstract
The Gaussian process (GP) is a widely used method for analyzing large-scale data sets, including spatio-temporal measurements of nonlinear processes that are now commonplace in the environmental sciences. Traditional implementations of GPs involve stationary kernels (also termed covariance functions) that limit their flexibility, and exact methods for inference that prevent application to data sets with more than about ten thousand points. Modern approaches to address stationarity assumptions generally fail to accommodate large data sets, while all attempts to address scalability focus on approximating the Gaussian likelihood, which can involve subjectivity and lead to inaccuracies. In this work, we explicitly derive an alternative kernel that can discover and encode both sparsity and nonstationarity. We embed the kernel within a fully Bayesian GP model and leverage high-performance computing resources to enable the analysis of massive data sets. We demonstrate the favorable performance of our novel kernel relative to existing exact and approximate GP methods across a variety of synthetic data examples. Furthermore, we conduct space-time prediction based on more than one million measurements of daily maximum temperature and verify that our results outperform state-of-the-art methods in the Earth sciences. More broadly, having access to exact GPs that use ultra-scalable, sparsity-discovering, nonstationary kernels allows GP methods to truly compete with a wide variety of machine learning methods.
@@ 1 Introduction
Gaussian processes (GPs) are a preeminent framework for stochastic function approximation, statistical modeling of real-world measurements, non-parametric and nonlinear regression within machine learning (ML), and surrogate modeling. GPs are analytically tractable and can be fully specified in terms of a prior mean function and a kernel, also known as a covariance function. Additionally, unlike most ML methods, GPs are “probabilistic” in the sense that they naturally include measures of uncertainty in predictions and, when implemented in a Bayesian setting, can furthermore account for and propagate uncertainty in their inferred hyperparameters. GPs have been broadly used across many areas of science, including statistical modeling of spatio-temporal data [1, 2], forward modeling [3, 4, 5], uncertainty quantification [6, 7], and autonomous experimentation [8, 9, 10, 11].
[…]
@@
While GPs are a widely-used method in many areas of science, they retain two properties that limit their broader application to modern, large-scale data sets, particularly in the age of artificial intelligence and machine learning. First, traditional implementations of GPs involve stationary kernels [12] that are inflexible for modeling real-world data [13]. The fact that standard stationary kernels are so widely used leads to the general impression that GPs themselves are inflexible, as opposed to a particular implementation of GPs being inflexible. Second, evaluating the Gaussian likelihood for $N$ data points scales with $\mathcal{O}(N^{2})$ in memory and $\mathcal{O}(N^{3})$ in computation, such that GPs are computationally infeasible for as few as ten thousand data points. Given the fact that modern data sets quite commonly involve millions or even billions of data points, in an ML context, this lack of scalability leads users to gravitate towards neural networks and deep learning methods that are easily applied to big data.
@@
The statistics and ML literature contains a wide variety of methods to address these two primary limitations of GPs. Much work has been done to derive and implement nonstationary kernels, wherein the dependence between two input locations cannot be written as a function of their separation distance [see, e.g., 14, 15, 16, 17, 18, 19]. However, with a few exceptions, these methods are generally difficult to implement for large data sets due to their large numbers of hyperparameters and the need for specialized algorithms for training. These methods can also lead to overfitting without suitable regularization techniques [20, 21, 22, 23, 24]. Regarding scalability, a variety of methods have been proposed that derive approximations to the GP likelihood, e.g., local GP experts [25], predictive process and low-rank methods [26, 27, 28], covariance tapering [29, 30], structure-exploiting methods [31], and Vecchia approximations [32, 33, 34]. We refer the interested reader to [35] for a direct comparison of many of these methods on a common data set. These so-called “approximate GP” methods are not, in general, agnostic to the underlying kernel, and theoretical properties are, in most cases, derived only for stationary kernels. Furthermore, these methods require subjective choices – e.g., Vecchia approximations necessitate specifying the number of neighbors to condition on, the ordering of points, and the selection of nearest neighbors – that may introduce sensitivities in a given analysis.
@@
A final limitation of most kernels used in the literature (both stationary and nonstationary) is that they cannot model covariances that are exactly zero. In almost all cases, as the separation between two points increases, a kernel will decay to (but not technically reach) zero, yielding dense covariance matrices. It is our hypothesis that most (if not all) real-world data sets implicitly include at least some degree of sparsity in the covariance, but typical classes of kernels cannot flexibly model both zero and nonzero covariances. Compactly-supported kernels [see, e.g., 36, 37, 38, 39, 40] can model both zero and nonzero covariances in a data-driven manner, albeit usually with stationary kernels. Approximate GP methods also “model” sparsity, albeit via ad hoc choices such as the ordering of points and neighbor selection. The combination of nonstationary and also compactly supported kernels would enable more appropriate modeling of real-world data and yield computational shortcuts for the matrix operations needed to calculate the Gaussian likelihood. In short, the density of the GP prior covariance is not fundamental to the method itself but a result of poorly chosen kernels. Instead of using these standard kernels and then retroactively working on scalability, we propose to give the GP the ability to discover sparsity during the model fitting process.
@@
Figure 1: Ground truth function over $[0,10]$ with 50 noisy measurements (panel a.); the synthetic function is both sparse and nonstationary. Panel (b) shows the resulting posterior mean function with 95% uncertainty intervals for six methods: a stationary Gaussian process with a squared-exponential (non-sparse) kernel, the Vecchia approximation [33], the predictive process [26], covariance tapering [29, 30], and our sparsity-discovering kernel combined with both stationary and nonstationary core kernels. Panel (c) shows posterior median covariance estimates (gray areas indicate where the kernel is identically zero). Our kernel has three clear benefits: first, it can discover sparsity in the data while using an exact Gaussian likelihood; second, it can infer the nonstationary structure in the data wherein diagonal elements of the covariance are larger at the endpoints of the domain; third, it outperforms exact and leading approximate methods in terms of root mean square error (RMSE) and continuous rank probability score (CRPS; lower scores for both metrics indicate a better fit).
@@
In this paper, we introduce a novel compactly supported kernel that can discover and encode sparsity, yielding ultra-flexible kernels that allow one to apply exact nonstationary GPs to arbitrarily large data sets. A demonstration of our proposed methodology is shown in Figure 1. Together with prior regularization, the induced sparsity also prevents overfitting and leads to a better interpretation of the data since it preserves the flexibility of a nonstationary kernel. We also propose a complete Bayesian stochastic model based on the new kernel that accounts for a generic prior mean function and measurement error and allows posterior prediction for unobserved input locations with uncertainty quantification. While high-performance computing (HPC) resources are needed to apply our method to truly massive data sets, we also show how the proposed kernel can be used to conduct fully Bayesian analysis for moderately sized data sets with limited computational resources, e.g., using a personal laptop. The efficacy of our novel kernel is demonstrated using both synthetic data examples and a real-world data set from the Earth sciences.
@@ 2 Sparsity-discovering nonstationary kernels
where $C_{\text{core}}$ is a positive definite kernel (stationary or nonstationary), $C_{\text{sparse}}$ is a sparsity-inducing kernel (defined below in Eq. 2), and ${\boldsymbol{\theta}}_{y}=({\boldsymbol{\theta}}_{\text{core}},{\boldsymbol{\theta}}_{\text{sparse}})$. While similar in appearance to covariance tapering [29, 30], our approach has important distinctions which will be discussed in the next section. So long as $C_{\text{core}}$ and $C_{\text{sparse}}$ are positive definite, their product will be positive definite [41], yielding a powerful framework for building flexible kernels.
@@ 2.1 Nonstationary kernels that discover sparsity
[…]
where $a\geq 0$ represents the amplitude, $b>0$ is a shape parameter (see Figure F.1 for a visualization of the effect of $b$), $\mathbf{h}\in\mathbb{R}^{d}$ is the centroid, and $r>0$ is the radius [42]. Note that $g({\bf x})$ is smooth and exactly zero for point-to-centroid distances greater than or equal to $r$, which is what enables the covariance matrix to be sparse. We combine these bump functions to yield a sparse kernel
@@
$f_{i}({\bf x})=\sum_{j=1}^{n_{2}}g_{ij}({\bf x}),\qquad i=1,\dots,n_{1},$
$f_{0}({\bf x},{\bf x}^{\prime};r_{0})$ is a compactly supported stationary kernel, and $s_{0}$ is a scalar parameter to adjust the relative magnitude of $f_{0}({\bf x},{\bf x};r_{0})$ and the amplitudes $\{a_{ij}\}$. We include an additive stationary kernel in Eq. 2 because there might otherwise be cases where $C({\bf x},{\bf x};{\boldsymbol{\theta}}_{\text{sparse}})=0$; specifying the stationary kernel to be compactly supported ensures that the resulting covariance matrix is still sparse. Among the broad range of possibilities for specifying a compactly supported $f_{0}$ [see, e.g., 37, 38, 39, 40], we opt to use members of the Wendland family of polynomials [36], e.g.,
[…]
@@
Before proceeding, we first verify that $C_{\text{sparse}}$ as defined in Equation 2 is a positive definite function and yields positive definite covariance matrices.
Proposition 1.
The kernel $C_{\text{sparse}}$ is strictly positive definite. Furthermore, $C_{y}$ is strictly positive definite whenever $C_{\text{core}}$ is.
The proof is provided in Appendix A.
As mentioned previously, using the product of covariances has previously been developed under the term “covariance tapering” [29, 30] and by now has almost twenty years of history. Indeed, without the additive cross-product term in Equation 2, the methodology would be essentially identical. However, the primary limitation of covariance tapering is its suboptimal performance when the length-scale of the compactly supported kernel (or “taper”, i.e., $r_{0}$) is small relative to the length-scale of the data. In other words, when large-scale or long-range correlations are present in the data, covariance tapering is not the best kernel choice for use within a GP. Such long-range correlations are widespread in data sets in the Earth sciences because of so-called “teleconnections” (defined as statistical relationships between weather patterns in distant locations), e.g., the El Niño/Southern Oscillation [43]. Importantly, in our formulation, use of the bump functions enables the kernel to capture long-range, off-diagonal correlations (see Figure 2) such that it maintains strong performance across a wide range of underlying data structures. We compare and contrast our method with covariance tapering directly in Section 4.2.
In order to show how the various components of $C_{\text{sparse}}$ in Eq. 2 combine to flexibly model sparsity in the overall kernel $C_{y}$, we present an illustrative one-dimensional example. We set $n_{1}=2$ and $n_{2}=3$ and take $\mathcal{X}=[0,1]$ with $N=100$ equally spaced location. Figure 2(a) shows the bump functions $g_{ij}(x)$ and their sums $f_{i}(x)$ using fixed values of the bump function hyperparameters. Starting with the leftmost column of Figure 2, the product of the sums-of-bump-functions $\sum_{i}f_{i}(x)f_{i}(x^{\prime})$ in the middle row reveals the effect of the bump functions: for a given $i$, the support (i.e., domain of nonzero values) of the bump functions define distance-unrelated parts of the domain that are correlated; furthermore, across the various $i$, a given point in the input space can have a nonzero correlation with different sets of points that does not depend on the distances from this point to these sets. For example, note that both because $f_{1}(0.17)>0$ and $f_{1}(0.7)>0$, $C_{\text{sparse}}(0.17,0.7)>0$ even though these points are relatively far apart. Furthermore, because $f_{2}(0.17)>0$ and $f_{2}(0.9)>0$ (but $f_{2}(0.7)=0$), we have $C_{\text{sparse}}(0.17,0.9)>0$ but $C_{\text{sparse}}(0.7,0.9)=0$. The radii $r_{ij}$ specify the size of distance-unrelated regions with nonzero correlations, while the amplitudes $a_{ij}$ and shapes $b_{ij}$ determine the relative magnitude of off-diagonal nonzero correlations and how sharply peaked the correlations are near the bump function centers (Figure 2b and c). Finally, increasing the radii in Figure 2(d) predictably decreases the overall sparsity of $C_{\text{sparse}}$ and (together with relatively small $b$) increases the areas of high correlations. The compactly supported stationary covariance $f_{0}(x,x^{\prime};r_{0}=0.15)$ serves to ensure that the diagonal elements of $C_{\text{sparse}}$ are nonzero, with $r_{0}$ determining the extent to which points very close together are correlated within the sparse kernel.
[…]
@@
Our specification of $C_{\text{sparse}}$ implies a nonstationary kernel, irrespective of the choice of $C_{\text{core}}$: note that we cannot write Eq. 2 as a function of the separating vector ${\bf x}-{\bf x}^{\prime}$ since the bump functions in Eq. 1 cannot be written as a function of ${\bf x}-{\bf x}^{\prime}$. Our approach explicitly models sparsity directly in the covariance matrix to capture the inherent sparsity of the dependence structure in real-world datasets; alternatively, approximate methods [e.g., inducing points; 44, 28] focus on approximating the GP with a subset of the data to achieve computational savings. In other words, both classes of methods leverage sparsity to achieve computational savings, but the sparsity implied by our method is data driven and describes the underlying structure in the data set of interest.